A bullet 10.1g leaves the barrel of a rifle with a speed of 504m/s. If the length of the barrel is 52.2cm determine the average force that accelerate the bullet.
1
Expert's answer
2021-09-29T09:54:53-0400

The average force is


"F=ma=m\frac{v^2}{2d}=0.0101·\frac{504^2}{2·0.522}=2457\text{ N}."